OpenStudy (anonymous):

Find the factors of 3a^2 + 14a − 24. (3a − 6)(a + 4) (3a − 4)(a + 6) (3a − 2)(a + 12) (3a + 2)(a − 12) i got the first option

OpenStudy (anonymous):

How did you get that?

OpenStudy (goldphenoix):

Hmm. Not quite, if you do the FOIL method on a, then it would give you something other than 3a^2 + 14a − 24.

OpenStudy (goldphenoix):

I'm 12, and I only learned the basic way of factoring a trinomial. So if it was me doing that problem, I would guess and check. Though, I'm sure other people knows how to do it, properly. :)

OpenStudy (anonymous):

if you need to know the true answer you must check

OpenStudy (anonymous):

14a

OpenStudy (anonymous):

\[3a ^{2} + 14a − 24. \] \[3a ^{2} +18a -4a − 24. \] \[3a(a+6) -4(a+6)\] (3a-4) (a+6)

OpenStudy (anonymous):

18a-4a=14a which one give you this solution?????

OpenStudy (anonymous):

I got B

OpenStudy (goldphenoix):

If you do the FOIL method on b, then that would certainly give you 3a^2 + 14a − 24. Good job, @FilthyMcNasty :)

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

@sylinan, you understand how i got B?

OpenStudy (anonymous):

Yes, No, Maybe so?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

yes (:

OpenStudy (goldphenoix):

Hmm. Teach me. c: I know that 18a - 4a = 14a. But how did you know which number to use?

OpenStudy (anonymous):

\[3a ^{2}+14a−24.\] (you multiply the 1st and Last terms) so 3a^2*-24 = -72a^2 - So u find 2 terms whose product is -72a^2 & add up to 14a

OpenStudy (goldphenoix):

Ah. I see. Thanks! :)

OpenStudy (anonymous):

3a²+14a−24. (you multiply the 1st and Last terms) so 3a²*-24 = -72a² - So u find 2 terms whose product is -72a² & add up to 14a

OpenStudy (anonymous):

If that's more legible , so yeah!